OpenStudy (anonymous):

Write an equation of the line that is parallel to the given line y = x + 3 and the given point (5,0).

OpenStudy (anonymous):

if you ask this questin on yahoo answers they might help you faster this is to hard for me :P

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

parallel lines have equal slopes for the line of the form y=mx+c m is the slope so the slope of y = x + 3 or y=1x+3 is..

OpenStudy (anonymous):

ok ths antonio

OpenStudy (anonymous):

no problem :))

OpenStudy (anonymous):

the equation of the line with slope m and passingh through (x1,y1) is given by y-y1= m(x-x1)

OpenStudy (anonymous):

can you proceed with the above information

OpenStudy (anonymous):

ok matricked thks a lot :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:) thks 4 tha medal

OpenStudy (anonymous):

alternatively eq of the line parallel to y=x+3 is y=x+k you may find k using the statemnet that the reqd line passes through (5,0)

OpenStudy (anonymous):

yw

OpenStudy (kaptain_mittens):

Sorry but I need help on a question I asked 40 minutes ago but no one has helped can either of you come check out my question and help me?

OpenStudy (anonymous):

@matricked is it y=1x+3

OpenStudy (anonymous):

is that tha answer

OpenStudy (anonymous):

pls i need 2 know tha answer 4 my test pls *-*

OpenStudy (anonymous):

can some1 help pls

OpenStudy (anonymous):

@paki

OpenStudy (paki):

matrickid has already explained...

OpenStudy (anonymous):

but is tha answer y=1x+3

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

is that rite paki

OpenStudy (paki):

use this.... "the equation of the line with slope m and passingh through (x1,y1) is given by y-y1= m(x-x1)"

OpenStudy (anonymous):

so tha answer is not rite

OpenStudy (anonymous):

is it y=mx

OpenStudy (anonymous):

i am so xD

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

sigh

OpenStudy (anonymous):

im tired nd cranky can som1 pls explain more

OpenStudy (paki):

@sahrya sorry, i am at office... have a look here... http://www.algebra.com/algebra/homework/Graphs.faq.question.388432.html

OpenStudy (anonymous):

ok thks baby :) call me sometime

OpenStudy (paki):

hhahah sure :)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

luv u <3 *kiss